Andre is waiting for the school bus. He knows that 2 out of 6 cars in his town are white. He wants to determine the probability that at least two white cars drive by in a row. He uses a number cube as a simulation. He lets the numbers 1 and 2 represent a white car and lets all other numbers represent a car that is not white. He rolls the number cube 6 times and records the outcomes in the table below.


1
Expert's answer
2021-04-15T06:56:44-0400

The possibility to roll 1 or 2 is "\frac{1}{3}". Thus, the possibility to roll 1 or 2 will be rolled two times in a raw = "(\frac{1}{3})^2" = "\frac{1}{9}"

Answer: "\frac{1}{9}."